# Stochastic Model Predictive Control of Charging Energy Hubs with Conformal Prediction

### v1

@@ V-B Control Performance
Figure 5: Energy hub operation during a one-day evaluation using the Stochastic MPC with a 96-step horizon window with a step resolution $\Delta T=0.25\,\mathrm{h}$. At each $k$, we present all scenarios obtained from the probabilistic 24 h-ahead predictions of $P_{\mathrm{ev}}$ and $P_{\mathrm{pv}}$, as the residual load ($\boldsymbol{P}_{\mathrm{ev},k}^{s}-\boldsymbol{P}_{\mathrm{pv},k}^{s}$) along with the observed residual load ($P_{\mathrm{ev},0|k}-P_{\mathrm{pv},0|k}$). We include the optimal trajectories obtained from the Stochastic MPC at each $k$ ($\boldsymbol{P}_{\mathrm{b},k}^{*s}$, $\boldsymbol{P}_{\mathrm{g},k}^{*s}$ and $\boldsymbol{E}_{\mathrm{b},k}^{*s}$ ). We also present, at each $k$, the observed $P_{\mathrm{g},0|k}$ and $E_{\mathrm{b},1|k}$ computed in the simulated environment of the energy hub, based on the applied action ($P_{\mathrm{b},0|k}^{*}$) and $P_{\mathrm{ev},0|k}-P_{\mathrm{pv},0|k}$.
In Fig. 5, we present the result of a one-day evaluation using the Stochastic MPC defined in Problem (1). Then, we compare the Stochastic MPC with the Recourse, the Deterministic and the Omniscient MPC during 280 days of evaluation. In Tab. IV we report the results using as a metric the observed daily net cost of operation. The Stochastic and Recourse MPC decrease the average cost of operation by 0.8%, when compared to the Deterministic Version. This also highlights that, in this case, using a point-estimate prediction with no measure of uncertainty (Deterministic MPC) yields acceptable performance. The slight performance improvement after accounting for variable uncertainties, does not come at the cost of critical computational time. In Tab. V, we show that Stochastic and Recourse MPC do not increase episodic computational time over 0.75 min, leaving enough time to compute optimal actions at every step $k$, by remaining far below $\Delta T$.
Table IV: Normalized average daily observed $C_{\mathrm{el}}$ [%] per season and per MPC version during a 280-day test. We normalized the values with the average daily observed $C_{\mathrm{el}}$ of the Omniscient MPC.
@@
Spring 100 115.75 115.44 115.45
Summer 100 108.14 107.45 107.45
Autumn 100 114.16 113.26 113.27
All 100 112.28 111.49 111.49

### v2

@@ V-B Control Performance
In Fig. 5, we present the result of a one-day evaluation using the Stochastic MPC defined in Problem (1). Then, we compare the Stochastic MPC with the Recourse, the Deterministic and the Omniscient MPC during 280 days of evaluation. In Tab. III and IV we report the results using as metrics the observed daily net cost of operation and the daily net $\mathrm{CO_{2}}$ emissions of the grid, respectively. The Stochastic and Recourse MPC decrease the average cost of operation by 0.9% and the average $\mathrm{CO_{2}}$ emissions by 0.3%, when compared to the Deterministic Version. This also highlights that, in this case, using a point-estimate prediction with no measure of uncertainty (Deterministic MPC) yields acceptable performance. The slight performance improvement after accounting for variable uncertainties, does not come at the cost of critical computational time. In Tab. V, we show that Stochastic and Recourse MPC do not increase episodic computational time over 0.75 min, leaving enough time to compute optimal actions at every step $k$, by remaining far below $\Delta T$.
Table III: Normalized average daily observed $C_{\mathrm{el}}$ [%] per season and per MPC version during a 280-day test. We normalized the values with respect to the Omniscient MPC.
@@
Spring 100 115.29 114.99 115.0
Summer 100 109.29 108.51 108.5
Autumn 100 116.41 115.37 115.37
All 100 113.65 112.76 112.77
Table IV: Normalized average daily observed $\mathrm{CO_{2}}$ emissions [%] per season and per MPC version during a 280-day test. We normalized the values with respect to the Omniscient MPC.
Season Omniscient Deterministic Stochastic Recourse
Winter 100 100.98 100.73 100.73
Spring 100 102.3 102.1 102.08
Summer 100 100.72 100.33 100.33
Autumn 100 101.75 101.54 101.54
All 100 101.4 101.13 101.13
@@
Figure 5: Energy hub operation during a one-day evaluation using the Stochastic MPC with a 96-step horizon window with a step resolution $\Delta T=0.25\,\mathrm{h}$. At each $k$, we present all scenarios obtained from the probabilistic 24 h-ahead predictions of $P_{\mathrm{ev}}$ and $P_{\mathrm{pv}}$, as the residual load ($\boldsymbol{P}_{\mathrm{ev},k}^{s}-\boldsymbol{P}_{\mathrm{pv},k}^{s}$) along with the observed residual load ($P_{\mathrm{ev},0|k}-P_{\mathrm{pv},0|k}$). We include the optimal trajectories obtained from the Stochastic MPC at each $k$ ($\boldsymbol{P}_{\mathrm{b},k}^{\star s}$, $\boldsymbol{P}_{\mathrm{g},k}^{\star s}$ and $\boldsymbol{E}_{\mathrm{b},k}^{\star s}$ ). We also present, at each $k$, the observed $P_{\mathrm{g},0|k}$ and $E_{\mathrm{b},1|k}$ computed in the simulated environment of the energy hub, based on the applied action ($P_{\mathrm{b},0|k}^{\star}$) and $P_{\mathrm{ev},0|k}-P_{\mathrm{pv},0|k}$.